OpenStudy (anonymous):

Solve for x, given the equation Square root of x minus 5 + 7 = 11

OpenStudy (anonymous):

\[\sqrt{x - 5} + 7 = 11\]

OpenStudy (anonymous):

x = 21, solution is extraneous x = 21, solution is not extraneous x = 81, solution is extraneous x = 81, solution is not extraneous

OpenStudy (nubeer):

1) First subtract 7 on both sides. 2) Then square both sides..

OpenStudy (nubeer):

3) Then add 5 on both sides.

OpenStudy (anonymous):

I subtract the other 7 from 5 or 11?

OpenStudy (anonymous):

11 right?

OpenStudy (nubeer):

subtract 7 from 11.. 11-7

OpenStudy (anonymous):

x = 9

OpenStudy (anonymous):

would the x be square root still?

OpenStudy (nubeer):

(x-5) ^0.5 = 11-7 (x-5)^0.5 = 4 now square both sides (x-5) = 4^2

OpenStudy (anonymous):

ohh yeah I forgot to square ughh sorry

OpenStudy (nubeer):

no problem.

OpenStudy (anonymous):

how do I square it?

OpenStudy (xapproachesinfinity):

square both side\(\Large \tt\color{blueviolet}{(\sqrt{x-5})^2=4^2\\ x-5=16}\)

OpenStudy (xapproachesinfinity):

solve for x now

OpenStudy (anonymous):

x - 25 = 16? @xapproachesinfinity

OpenStudy (xapproachesinfinity):

why x-25? how did you come with that

OpenStudy (xapproachesinfinity):

you just need to add 5 to both side this linear equation now?

OpenStudy (anonymous):

ohh I hasd ssquare rooted it

OpenStudy (xapproachesinfinity):

x-5=16 5 5 _______________ x=21

OpenStudy (xapproachesinfinity):

No! square root with the square power cancel

OpenStudy (anonymous):

ugh! Sorry, now how do I find out if its extraneous or not?

OpenStudy (xapproachesinfinity):

i'm not sure i understand what you meant? you were looking for x and you got the solution?

OpenStudy (anonymous):

I know but the choices are x = 21 is extraneous and x = 21, solution is not extraneous @xapproachesinfinity

OpenStudy (nubeer):

It is ..

OpenStudy (anonymous):

Its wrong

OpenStudy (xapproachesinfinity):

what does that mean by the way never heard such a thing ??

OpenStudy (nubeer):

@xapproachesinfinity google it :P :P

OpenStudy (xapproachesinfinity):

now i got it^_^ i google it actually okay you answer is it not extraneous the reason is 21 was included in the domain of the equation to know if this is true do this x-5>=0 x>5 so all real number that are greater than or equal to 5 therefore 21 was included this thing is simple when you solve watch out for the domain especially the rational expression logarithm and such

OpenStudy (xapproachesinfinity):

@_angeliquexoxo it is good to know how to do this for next time so watch the domain^_^

OpenStudy (phi):

fyi, an extraneous solution is an "answer" you get after doing the algebra (correctly, of course), but when you put the answer into the original problem, *it does not work* however, in this case, if we put 21 into your problem (i.e. replace x with 21) and simplify we get \[ \sqrt{21-5}+7 \\ \sqrt{16}+7 \\ 4+7\\11 \] which matches the right-hand side. in other words, 21 "works" so it is *not extraneous*

OpenStudy (anonymous):

oh okay thankyou for the explanation! (: @phi

OpenStudy (phi):

the English (not math) definition of extraneous is Inessential or unrelated to the topic or matter at hand; irrelevant. which might help you to remember what the "math" definition means